3. If you are taking a random sample from a normally distributed 5-12, determine if the following outcomes are a likely or unlike a. A sample mean greater than 106 for a sample of n4. b. A sample mean greater than 106 for a sample of n -36 population with ?-100 and ly event. (4 pts)
Answer #1

Here we have given that

\mu = population mean =100

\sigma= population standard deviation=12

(A)

Now we want to calculate the probability of sample mean is greater than 106

for n=sample size =4

i.e P( \bar x > 106)

for finding this probability 1st we need to find the Zscore

Zscore= \frac{\bar x- \mu } { \frac{\sigma}{\sqrt(n)}}

=\frac{106 - 100 } { \frac{12}{\sqrt(4)}}

=1.00

i.e we get

P(Z > 1.00 ) = 1- P( Z < 1.00)

= 1 - 0.8413 using z standard normal table

=0.1587

we get

the probability of sample mean is greater than 106 =0.1587

i.e P( \bar x > 106) =0.1587

(B)

Now we want to calculate the probability of sample mean is greater than 106

for n=sample size =36

i.e P( \bar x > 106)

for finding this probability 1st we need to find the Zscore

Zscore= \frac{\bar x- \mu } { \frac{\sigma}{\sqrt(n)}}

=\frac{106 - 100 } { \frac{12}{\sqrt(36)}}

=3.00

i.e we get

P(Z > 3.00 ) = 1- P( Z < 3.00)

= 1 - 0.9987 using z standard normal table

=0.0013

we get

the probability of sample mean is greater than 106 for n=36 is 0.1587

i.e P( \bar x > 106) =0.0.0013